OpenStudy (anonymous):

What is the best way to classify each equation? Column A Column B 1. 2. 3. 4. A. identity B. contradiction C. neither

OpenStudy (anonymous):

oh shoot hold on

OpenStudy (anonymous):

1. 22-3x+7x=4(x+5)

OpenStudy (anonymous):

2. 6(2x-3)=3(3x-5)

OpenStudy (anonymous):

22 - 3x + 7x = 4(x + 5) 22 + 4x = 4(x + 5) 22 + 4x = 4x + 20 22 - 20 + 4x = 4x + 20 - 20 2 + 4x - 4x = 4x - 4x 2 = 0

OpenStudy (anonymous):

3. 6x-14=2(3x-7)

OpenStudy (anonymous):

4. 6x+3(x-4)=8(x-3)

OpenStudy (anonymous):

6(2x-3)=3(3x-5) 12x - 18 = 3(3x - 5) 12x - 18 = 9x - 15 12x - 18 + 18 = 9x -15 + 18 12x - 9x = 9x - 9x + 3 3x = 3 x = 1

OpenStudy (anonymous):

6x-14=2(3x-7) 6x - 14 = 6x - 14

OpenStudy (anonymous):

6x+3(x-4)=8(x-3) 6x + 3x - 12 = 8x - 24 9x - 12 = 8x - 24 9x - 12 + 24 = 8x - 24 + 24 9x + 12 = 8x 9x - 9x + 12 = 8x - 9x 12 = -x -12 = x

OpenStudy (anonymous):

so 1. b 2. c 3.a and 4. a?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

k thanks I will tell u what score I got

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

60% I failed :,(

OpenStudy (anonymous):

I will tell u what I got wrong brb

OpenStudy (anonymous):

That's strange, did it only consist of the questions you asked me?

OpenStudy (anonymous):

What did you get wrong?

OpenStudy (anonymous):

wait like 15 min.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes all the questions u saw were on the test

OpenStudy (anonymous):

Were any of the ones I helped you with wrong?

OpenStudy (anonymous):

hang on my mom has to put in the password so I can see what I got wrong

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

ok im back

OpenStudy (anonymous):

Okay, what did you get wrong?

OpenStudy (anonymous):

so we got the first one and the last one wrong

OpenStudy (anonymous):

Hmm. Let me check the first one again

OpenStudy (anonymous):

its was supposed to be one solution

OpenStudy (anonymous):

and on number five the fourth one was supposed to be c one solution

OpenStudy (anonymous):

@bloonscat where did u go